ADNAN MUHAMMED

Grade 12,

i have 2 questions --- The Condition that the roots of the eqn ax 2 + bx + c = 0 such that one root is n times the other is : a) na 2 = bc(n+1) 2 b) nb 2 = ca (n+1) 2 c) nc 2 =ab(n+1) 2 2. the root of the eqn ax 2 +bx+c = 0 are in the ratio m:n, then a )mna 2 =(m+n)c 2 b)mnb 2 =(m+n)ac c)mnb 2 =(m+n)2ac

i have 2 questions  ---
 
  1. The Condition that the roots of the eqn ax+ bx + c = 0 such that one root is n times the other is : 
              a) na2= bc(n+1)
               b) nb2= ca (n+1)2
              c) nc2=ab(n+1)2        
       
       2. the root of the eqn ax2+bx+c = 0 are in the ratio m:n, then
 
          a)mna2=(m+n)c2
          b)mnb2=(m+n)ac
          c)mnb2=(m+n)2ac

Grade:10

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
1)let the roots be x,xn
so x+xn=(-b/a) and x2n=c/a
x=(-b/a(n+1)).........(1) and also x2=(c/an).......(2)
substituting 1 in 2 we get
nb2= ca (n+1)2
2)let (x/y)=(m/n)
So km,kn are roots of ax2+bx+c=0
so k(m+n)=-b/a and
k2mn=c/a
so (m+n)2/mn=(b/a)2+(c/a)
So mnb2=(m+n)2ac